# Performance Assessment of Lidar Odometry Frameworks: A Case Study at the Australian Botanic Garden Mount Annan

### v1

@@ Abstract
Autonomous vehicles are being tested in diverse environments worldwide. However, a notable gap exists in evaluating datasets representing natural, unstructured environments such as forests or gardens. To address this, we present a study on localisation at the Australian Botanic Garden Mount Annan. This area encompasses open grassy areas, paved pathways, and densely vegetated sections with trees and other objects. The dataset was recorded using a 128-beam LiDAR sensor and GPS and IMU readings to track the ego-vehicle. This paper evaluates the performance of two state-of-the-art LiDAR-inertial odometry frameworks, COIN-LIO and LIO-SAM, on this dataset. We analyse trajectory estimates in both horizontal and vertical dimensions and assess relative translation and yaw errors over varying distances. Our findings reveal that while both frameworks perform adequately in the vertical plane, COIN-LIO demonstrates superior accuracy in the horizontal plane, particularly over extended trajectories. In contrast, LIO-SAM shows increased drift and yaw errors over longer distances.
@@ 1 Introduction
Despite advancements in these frameworks, there remains a need for comprehensive evaluations in diverse and challenging environments. The Australian Botanic Garden Mount Annan in Fig. 1. provides an ideal setting for such an evaluation due to its complex terrain and rich vegetation, which pose significant challenges for SLAM systems. The presence of a looped path and varying environmental conditions make it suitable for testing the robustness and accuracy of loop closure mechanisms.
This paper aims to evaluate and compare the performance of COIN-LIO and LIO-SAM when applied to the collected dataset. We focus on their ability to maintain high-accuracy localisation without loop closure.
The paper is organised as follows: Section 2 reviews related work in LiDAR-inertial odometry and loop closure techniques. Section 3 describes the methodology, including dataset details, implementation of the frameworks, and evaluation metrics used. Section 4 presents the results and analysis of the trajectory estimates, position drift, and error metrics. Section 5 discusses the findings, explores potential reasons for the observed performance differences, and considers implications for future work. Finally, Section 6 concludes the paper and outlines possible directions for further research.
2 Literature Review
@@ 2.1 LiDAR-Inertial Odometry
LiDAR-inertial odometry combines LiDAR data with inertial measurements to provide continuous and accurate localization. The IMU provides high-frequency motion estimates, helping to bridge gaps between LiDAR scans, especially in dynamic or geometrically challenging environments, such as tunnels, flat fields, or dense forests. Several advanced frameworks have been developed to enhance this synergy and improve system robustness, particularly with Ouster sensors.
LEGO-LOAM[3] is a foundational framework that optimizes the structure of traditional LOAM by reducing computational complexity while maintaining real-time performance. However, it does not natively integrate IMU data, which can limit its performance in high-speed, dynamic environments.
LIO-SAM[4], on the other hand, directly addresses this by tightly coupling LiDAR and IMU data through a factor graph approach, making it well-suited for aggressive motion scenarios and reducing drift. This integration significantly improves its performance with Ouster sensors in particular.
FAST-LIO[5] builds on this by providing a lightweight and computationally efficient real-time odometry solution. Its architecture is ideal for high-speed applications and performs exceptionally well with multi-beam LiDARs like those from Ouster, enabling precise localization even in challenging environments.
Moreover, Imaging Lidar Place Recognition[6] focuses on robust place recognition and loop closure, which is essential for ensuring global consistency in long-duration SLAM systems. This framework complements odometry methods and performs well with multi-beam imaging lidars, enhancing the overall robustness in challenging conditions.
Lastly, a notable recent development is COIN-LIO, an innovative framework introduced in 2024. Built on the efficient FAST-LIO2[7] architecture, it improves the robustness of LiDAR-inertial odometry in geometrically degenerate environments such as tunnels and flat fields. By projecting LiDAR intensity returns, similarly as in the Imaging Lidar Place Recognition, into an image and introducing a novel image processing pipeline, it produces filtered images with enhanced brightness consistency(Fig. 2), especially while being highly compatible with Ouster sensors.
@@
We set a comparative table (Fig. 3) to distinguish between the characteristics of each of the afore-mentioned frameworks, prior to starting the experimental setup.
Figure 3: Table showing promising selection criteria of LIO-SAM and COIN-LIO over other frameworks
@@ 2.2 Trajectory Evaluation Tools
Evaluating the performance of LiDAR-inertial odometry systems necessitates precise tools for comparing estimated trajectories against ground truth data. Two prominent tools in this domain are the RPG Trajectory Evaluation Toolbox[8] and EVO (Evolution)[9].
The RPG Trajectory Evaluation Toolbox, developed by the Robotics and Perception Group at the University of Zurich, offers a comprehensive framework for trajectory analysis. It provides detailed metrics such as Absolute Trajectory Error (ATE) and Relative Pose Error (RPE), which are essential for assessing both global consistency and local accuracy of the trajectories. The toolbox is known for its robustness and ease of integration with various datasets and systems, making it a preferred choice for researchers focusing on high-precision evaluations.
EVO is another widely used tool for the evaluation of odometry and SLAM algorithms. It supports multiple data formats and offers visualization capabilities alongside quantitative metrics. EVO is particularly user-friendly and is compatible with common datasets like TUM RGB-D[10] and KITTI[11], allowing for flexible application across different evaluation scenarios.
In addition to these, other tools such as tum-eval and KITTI odometry evaluation toolkit exist, but they are more specialized towards their respective datasets.
In the context of this study, the RPG Trajectory Evaluation Toolbox was selected over EVO due to its advanced analytical capabilities and better alignment with the evaluation criteria of our LiDAR-inertial odometry systems. The RPG toolbox’s detailed error metrics and seamless integration with our experimental setup provided more insightful results, facilitating a deeper understanding of the system’s performance in complex environments.
@@ 2.3 Loop Closure Techniques
Recent approaches have utilized learning-based methods to improve loop closure detection. For instance, SegMatch uses segment-based place recognition to identify loop closures in 3D point clouds[9]. Others have incorporated semantic information to enhance matching robustness in dynamic environments.
@@
2.4 Challenges in Complex Environments
Navigating complex environments like urban areas or natural settings with dense vegetation poses significant challenges for LiDAR-inertial odometry systems. Factors such as dynamic obstacles, repetitive structures, and occlusions can affect the accuracy of scan matching and loop closure detection.
Environmental conditions, such as varying lighting and weather, can also impact sensor performance. The integration of robust outlier rejection methods and adaptive algorithms is essential for maintaining localization accuracy under these conditions.
@@ 2.5 Evaluation Metrics
Evaluating SLAM systems requires comprehensive metrics that capture both local and global accuracy. Common metrics include Absolute Trajectory Error (ATE) and Relative Pose Error (RPE) [10]. ATE measures the difference between the estimated trajectory and the ground truth, providing insight into the overall accuracy. RPE assesses the drift over time by comparing the relative motion between successive poses.
Other considerations include computational efficiency, real-time performance, and the ability to handle loop closures effectively. Benchmark datasets and standardized evaluation protocols are important for comparing different SLAM systems objectively.
@@ 3.1 Dataset Description
The dataset used for this study was collected at the Australian Botanic Garden Mount Annan, a large public garden known for its diverse flora, intricate pathways, and varied terrain. It is 140 seconds long, is of a 9.7GBs of size, and was recorded on the 28th of March 2024. As of that date, the environment (Fig. 5) included open grassy areas, dense vegetation, water features, and road sign features and a moving car at one point of the dataset.
The dataset comprises LiDAR point clouds and IMU data collected along a predefined looped trajectory covering approximately 890 meters. The trajectory includes sections with tight turns, elevation changes, and areas with limited features. Ground truth data was obtained using a NovAtel PwrPak7D-E1 GNSS receiver, ensuring centimeter-level accuracy.
@@
LIO-SAM’s IMU pre-integration parameters in its configuration file were configured with respect to the OS1-128 sensor we used for the bag recording. The framework’s loop closure detection was enabled, with parameters adjusted to accommodate the larger scale of the trajectory. The voxel grid filtering and scan matching parameters were set to values suitable for outdoor environments with varying feature densities.
@@ 4.1 Absolute Trajectory Error (ATE)
ATE measures the global accuracy of the estimated trajectory by comparing each estimated point to the corresponding groundtruth point after aligning the two trajectories. This metric is crucial to understanding how well the estimated trajectory follows the groundtruth over the entire duration.
@@
  • Alignment: The estimated trajectory is first aligned to the groundtruth using a rigid-body or similarity transformation. This process ensures that the starting points of the trajectories are in the same reference frame. Alignment is typically based on minimizing the positional error across the entire trajectory, often starting with the first state.
@@
    Where $p_{i}$ represents the position at the $i$-th timestamp.
@@ 4.2 Relative Error (RE)
RE focuses on the local accuracy of the trajectory, assessing how well the estimated trajectory maintains the relative positions and orientations between pairs of points along the path. Unlike ATE, RE is less sensitive to global drift but instead measures the trajectory’s consistency.
@@ 4.3 Role of Starting Point in Evaluation
For both ATE and RE, the starting point plays a crucial role. In ATE, the starting point is typically used to align the entire trajectory, ensuring that the initial positions and orientations are consistent between the estimated and groundtruth trajectories. In RE, each sub-trajectory’s starting point is aligned to the groundtruth, allowing for an accurate assessment of the relative error at the second point in the pair.
@@ 4.4 Experimental Procedure
The experimental procedure, which relied on the well renowned trajectory evaluator RPG-Trajectory [11], involved the following steps:
@@ 5.1 Trajectory Analysis on the x-y Plane
For LIO-SAM, the initial portion of the trajectory (Fig. 7) aligned well with the ground truth. However, significant deviations began to accumulate after approximately 700 meters into the trajectory. In the final quarter of the loop, the estimated path diverged noticeably from the ground truth, particularly in the x-y plane.
@@ 5.3 Position Drift Evaluation
The position drift over the trajectory length was plotted for the x, y, and z coordinates. For COIN-LIO, the drift remained within ±1 meter for the x and y axes and within ±0.8 meters for the z-axis. The drift patterns showed small oscillations but no significant accumulations.
@@
In contrast, LIO-SAM exhibited increasing drift in the x and y axes after 700 meters. The drift exceeded -4 meters in both axes by the end of the trajectory. The z-axis drift remained within ±1.5 meters, indicating that vertical estimation was less affected.
@@ 5.4 Translation and Yaw Error Analysis for LIO-SAM
Figure 11 illustrates various errors associated with the LIO-SAM system, providing insights into its accuracy across different traveled distances.
Translation Error as a Percentage: Figure 11(a) shows the translation error as a percentage of the distance traveled for LIO-SAM. At shorter distances (e.g., 95.22 meters), the relative error fluctuates between 4% and 6%, indicating variability in system performance. However, as the distance increases, the error stabilizes between 2% and 4%, suggesting that while absolute translation errors increase, the system maintains consistent proportional accuracy over longer distances.
@@
Overall Conclusion: Together, these subfigures demonstrate how LIO-SAM’s performance degrades over longer trajectories. While the relative translation error stabilizes as a percentage of distance, the absolute translation and yaw errors increase significantly with distance. This trend suggests that although LIO-SAM maintains proportional accuracy for relative distances, its absolute positional and angular errors grow, impacting precision over extended distances. Applications requiring high accuracy over long trajectories should account for this increase in error.
@@ 5.5 Translation and Yaw Error Analysis for COIN-LIO
Figure 12(a) shows the absolute translation error for COIN-LIO across distances. At shorter distances, the error is around 5 meters, increasing to 15 meters at 476.13 meters. The widening of the box plots suggests increasing error variability. Compared to LIO-SAM, COIN-LIO’s errors are generally lower, indicating better accuracy.
@@
Figure 12(c) shows yaw error. At shorter distances, it ranges from 0.2 to 0.4 degrees, increasing to 0.8 degrees at 476.13 meters. Despite this, COIN-LIO maintains better angular accuracy than LIO-SAM, demonstrating greater stability over longer distances.
@@
Figures 11 and 12 collectively show that COIN-LIO delivers better performance than LIO-SAM in both absolute and relative translation errors, particularly over longer distances. While the yaw error increases with distance, COIN-LIO’s angular precision remains robust. Overall, COIN-LIO demonstrates consistent and proportional accuracy, making it a reliable option for applications requiring precise positioning and orientation, even over longer trajectories.
@@ 5.6 Statistical Analysis
Statistical tests were conducted to assess the significance of the differences observed in translation and yaw errors.
@@
A paired t-test on the relative translation errors indicated that COIN-LIO consistently outperformed LIO-SAM across all distances (p ¡ 0.05). The mean relative translation error for COIN-LIO was 3.5 percents, compared to 4.5 percents for LIO-SAM.

### v2

@@ Abstract
Autonomous vehicles are being tested in diverse environments around the world. However, there is a notable gap in the evaluation of datasets representing natural unstructured environments such as forests or gardens. To address this, we present a study on localization at the Mount Annan Botanic Garden. This area encompasses open grassy areas, paved pathways, and densely vegetated sections with trees and other objects. The data set was recorded using a 128-beam LiDAR sensor and GPS and IMU readings to track the ego vehicle. This paper evaluates the performance of two state-of-the-art LiDAR-inertial odometry frameworks, COIN-LIO and LIO-SAM, on this dataset. We analyze trajectory estimates in both horizontal and vertical dimensions and assess relative translation and yaw errors over varying distances. Our findings reveal that while both frameworks perform adequately in the vertical plane, COIN-LIO demonstrates superior accuracy in the horizontal plane, particularly over extended trajectories. In contrast, LIO-SAM shows an increase in drift and yaw errors over longer distances.
@@ 1 Introduction
Despite advances in these frameworks, there remains a need for comprehensive evaluations in diverse and challenging environments. The Mount Annan National Botanic Garden in Fig. 1 provides an ideal setting for such an evaluation due to its complex terrain and rich vegetation, which pose significant challenges for SLAM systems. The presence of a loop path and varying environmental conditions make it suitable for testing the robustness and accuracy of loop closure mechanisms.
Other public datasets such as WildPlaces[3], is intended for use in lidar place recognition in unstructured natural environments, and is collected with a handheld sensor payload. MulRan[4] is another large-scale data set that captures mixed urban and suburban sequences taken with vehicle mounted systems in various terrains.
The Mount Annan Australian Botanic Garden, depicted in Fig. 1, presents an ideal setting in its turn due to its looped path and densely vegetated roadways. Through this vehicle-mounted system, we are complementing gaps that current datasets do not address.
This paper aims to evaluate and compare the performance of COIN-LIO and LIO-SAM when applied to the collected dataset.
Criteria LIO-SAM COIN-LIO LEGO-LOAM Imaging Lidar Place Localization FAST-LIO
Real-time Performance Variable
Tightly-coupled IMU Integration Scan-Context Compatible
Loop Closure Scan-Context Compatible Scan-Context Compatible Scan-Context Compatible
Computational Efficiency Moderate High High Moderate High
Mapping Capability Low Limited
Robustness to Dynamic Environments Moderate High Low High Moderate
Figure 2: Comparison table of Lidar Inertial Odometry Methods
2 Related Work
@@ 2.1 LiDAR-Inertial Odometry
LiDAR-inertial odometry combines LiDAR data with inertial measurements to provide continuous and accurate localization. The IMU provides high-frequency motion estimates, helping bridge gaps between LiDAR scans, especially in dynamic or geometrically challenging environments, such as tunnels, flat fields, or dense forests. Several advanced frameworks have been developed to enhance this synergy and improve the robustness of the system, particularly with Ouster sensors.
LEGO-LOAM[5] is a foundational framework that optimizes the structure of traditional LOAM by reducing computational complexity while maintaining real-time performance. However, it does not natively integrate IMU data, which can limit its performance in high-speed dynamic environments.
LIO-SAM, on the other hand, directly addresses this by tightly coupling LiDAR and IMU data through a factor graph approach, making it well suited for aggressive motion scenarios and reducing drift. This integration significantly improves its performance with Ouster sensors in particular.
Moreover, Imaging Lidar Place Recognition[6] focuses on robust place recognition and loop closure, which is essential to ensure global consistency in long-duration SLAM systems. This framework complements odometry methods and performs well with multibeam imaging lidars, enhancing the overall robustness in challenging conditions.
Lastly, a notable recent development is COIN-LIO, an innovative framework introduced in 2024. Built on the efficient FAST-LIO2[7] architecture, it improves the robustness of LiDAR-inertial odometry in geometrically degenerate environments such as tunnels and flat fields. By projecting LiDAR intensity returns into an image and introducing a novel image processing pipeline, it produces filtered images with enhanced brightness consistency(Fig. 3), especially while being highly compatible with Ouster sensors. COIN-LIO further leverages intensity as an additional modality by employing a new feature selection scheme that detects uninformative directions in point cloud registration and explicitly selects patches with complementary image information. Photometric error minimization in these image patches is then fused with inertial measurements and point-to-plane registration in an iterated Extended Kalman Filter.
@@
We set a comparative table (Fig. 2) to distinguish between the characteristics of each of the frameworks mentioned above, prior to starting the experimental setup.
@@ 2.2 Trajectory Evaluation Tools
Evaluating the performance of LiDAR-inertial odometry systems requires precise tools for comparing estimated trajectories against ground-truth data. Two prominent tools in this domain are the RPG Trajectory Evaluation Toolbox[8] and EVO (Evolution)[9].
The RPG Trajectory Evaluation Toolbox, developed by the Robotics and Perception Group at the University of Zurich, offers a comprehensive framework for trajectory analysis. Provide detailed metrics such as the Absolute Trajectory Error (ATE) and Relative Pose Error (RPE), which are essential for assessing both global consistency and local accuracy of the trajectories. The toolbox is known for its robustness and ease of integration with various datasets and systems, making it a preferred choice for researchers who focus on high-precision evaluations.
EVO is another widely used tool to evaluate odometry and SLAM algorithms. Supports multiple data formats and offers visualization capabilities in addition to quantitative metrics. EVO is particularly user-friendly and compatible with common data sets such as TUM RGB-D [10] and KITTI [11], allowing flexible application in different evaluation scenarios.
In addition to these, other tools such as tum-eval and KITTI odometry evaluation toolkit are more specialized for their respective datasets.
In the context of this study, the RPG Trajectory Evaluation Toolbox was selected over EVO due to its advanced analytical capabilities and better alignment with the evaluation criteria of our LiDAR-inertial odometry systems. The detailed error metrics in the RPG toolbox and seamless integration with our experimental setup provided more insightful results, facilitating a deeper understanding of the system’s performance.
@@ 2.3 Loop Closure Techniques
Recent approaches have used learning-based methods to improve loop closure detection. For example, SegMatch uses segment-based place recognition to identify loop closures in 3D point clouds [9]. Others have incorporated semantic information to enhance matching robustness in dynamic environments.
@@ 2.4 Evaluation Metrics
Evaluating SLAM systems requires comprehensive metrics that capture both local and global accuracy. Common metrics include the absolute trajectory error and the relative position error. ATE measures the difference between the estimated trajectory and ground truth, providing insight into the overall accuracy. RPE assesses the drift over time by comparing the relative motion between successive poses.
Other considerations include computational efficiency, real-time performance, and the ability to handle loop closures effectively. Benchmark data sets and standardized evaluation protocols are important to compare different SLAM systems objectively.
@@ 3.1 Dataset Description
The data set used for this study was collected at Mount Annan, Mount Annan Botanic Garden in Australia, a large public garden known for its diverse flora, intricate pathways and varied terrain. It is 140 s long, is 9.7GB in size, and was recorded on 28 March 2024. At that date, the environment (Fig. 5) included open grassy areas, dense vegetation, water features and road sign features, and a moving car at one point in the data set.
The data set comprises LiDAR point clouds and IMU data collected along a predefined looped trajectory that covers approximately 890 meters. The trajectory includes sections with tight turns, elevation changes, and areas with limited features. Ground truth data was obtained using a NovAtel PwrPak7D-E1 GNSS receiver, ensuring centimeter-level accuracy.
@@
LIO-SAM’s IMU pre-integration parameters in its configuration file were configured with respect to the OS1-128 sensor we used for the bag recording. The framework’s loop closure detection was enabled, with parameters adjusted to accommodate the larger scale of the trajectory. The filtering parameters of the voxel grid and the scanning matching were set to suitable values for outdoor environments with varying density of features.
@@ 4.1 Absolute Trajectory Error (ATE)
ATE measures the global accuracy of the estimated trajectory by comparing each estimated point with the corresponding ground truth point after aligning the two trajectories. This metric is crucial to understanding how well the estimated trajectory follows the ground truth throughout the duration.
@@
  • Alignment: The estimated trajectory is first aligned to the groundtruth using a rigid-body or similarity transformation. This process ensures that the starting points of the trajectories are in the same reference frame. Alignment is typically based on minimizing positional error throughout the trajectory, often starting with the first state.
@@
    Here, $p_{i}$ represents the position on the $i$-th timestamp.
@@ 4.2 Relative Error (RE)
RE focuses on the local precision of the trajectory, evaluating how well the estimated trajectory maintains the relative positions and orientations between pairs of points along the path. Unlike ATE, RE is less sensitive to global drift, but instead measures the consistency of the trajectory.
@@ 4.3 Role of Starting Point in Evaluation
For both ATE and RE, the starting point plays a crucial role. In ATE, the starting point is typically used to align the entire trajectory, ensuring that initial positions and orientations are consistent between the estimated and groundtruth trajectories. In RE, each subtrajectory’s starting point is aligned to the groundtruth, allowing for an accurate assessment of the relative error at the second point in the pair.
@@ 4.4 Experimental Procedure
The experimental procedure, which relied on the well-known trajectory evaluator RPG-Trajectory [11], involved the following steps:
@@ 5.1 Trajectory Analysis on the x-y Plane
For LIO-SAM, the initial portion of the trajectory (Fig. 7) aligned well with the ground truth. However, significant deviations began to accumulate after approximately 700 meters into the trajectory. In the final quarter of the loop, the estimated path diverged noticeably from the ground truth, particularly in the xy plane.
@@ 5.3 Position Drift Evaluation
The position drift over the length of the trajectory was plotted for the x, y, and z coordinates. For COIN-LIO, the drift remained within ±1 meter for the x and y axes and within ±0.8 meters for the z axis. The drift patterns showed small oscillations but no significant accumulations.
@@
In contrast, LIO-SAM exhibited an increasing drift in the x and y axes after 700 meters. The drift exceeded -4 meters in both axes at the end of the trajectory. The z-axis drift remained within ±1.5 meters, indicating that vertical estimation was less affected.
@@ 5.4 Translation and Yaw Error Analysis for LIO-SAM
Figure 11 illustrates various errors associated with the LIO-SAM system, providing information on its accuracy at different distances traveled.
Translation Error as a Percentage: Figure 11(a) shows the translation error as a percentage of the distance traveled for LIO-SAM. At shorter distances (e.g. 95.22 meters), the relative error fluctuates between 4% and 6%, indicating variability in system performance. However, as the distance increases, the error stabilizes between 2% and 4%, suggesting that while absolute translation errors increase, the system maintains a consistent proportional accuracy over longer distances.
@@
Overall Conclusion: Together, these subfigures demonstrate how the performance of LIO-SAM degrades over longer trajectories. Although the relative translation error stabilizes as a percentage of distance, the absolute translation and yaw errors increase significantly with distance. This trend suggests that although LIO-SAM maintains proportional accuracy for relative distances, its absolute positional and angular errors grow, impacting precision over extended distances. Applications requiring high accuracy over long trajectories should account for this increase in error.
@@ 5.5 Translation and Yaw Error Analysis for COIN-LIO
Figure 12(a) shows the absolute translation error for COIN-LIO across distances. At shorter distances, the error is around 5 meters, increasing to 15 meters at 476.13 meters. The widening of the box plots suggests increased error variability. Compared to LIO-SAM, the COIN-LIO errors are generally lower, indicating better accuracy.
@@
Figure 12(c) shows the yaw error. At shorter distances, it ranges from 0.2 to 0.4 degrees, increasing to 0.8 degrees at 476.13 meters. Despite this, COIN-LIO maintains better angular accuracy than LIO-SAM, demonstrating greater stability over longer distances.
@@
Figures 11 and 12 collectively show that COIN-LIO delivers better performance than LIO-SAM in absolute and relative translation errors, particularly over longer distances. Although the yaw error increases with distance, COIN-LIO’s angular precision remains robust. Overall, COIN-LIO demonstrates consistent and proportional accuracy, making it a reliable option for applications that require precise positioning and orientation, even over longer trajectories.
@@ 5.6 Statistical Analysis
Statistical tests were performed to assess the significance of the differences observed in translation and yaw errors.
@@
A paired t-test on relative translation errors indicated that COIN-LIO consistently outperformed LIO-SAM at all distances (p ¡ 0.05). The mean relative translation error for COIN-LIO was 3.5 percents, compared to 4.5 percents for LIO-SAM.